OpenStudy (anonymous):

A rectangle has vertices whose coordinates are A(1, 4), B(4, -1), C(-1, -4), and D(-4, 1). Find the midpoints of the diagonals of the rectangle. Midpoint of AC = (,) Midpoint of BD = (,)

jimthompson5910 (jim_thompson5910):

Midpoint of AC = ( (x1+x2)/2 , (y1+y2)/2 ) Midpoint of AC = ( (1+(-1))/2 , (4+(-4))/2 ) Midpoint of AC = ( 0/2 , 0/2 ) Midpoint of AC = ( 0, 0 ) Follow the same steps to find the midpoint of BD

OpenStudy (anonymous):

thanks soo much!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

is BD=0,0 too?

jimthompson5910 (jim_thompson5910):

yes it is, very nice work

OpenStudy (anonymous):

YAY:)

jimthompson5910 (jim_thompson5910):

lol glad you know what you're doing here